OpenStudy (anonymous):

What is the simplified form of the following expression? 6c^2+2.5-d+2c^2-3d A. 4c^2 - 1.5d B. 6c^2 + 0.5d C. 8c^2 - 1.5d D. 8c^2 + 1.5d

OpenStudy (anonymous):

@TheEdwardsFamily

OpenStudy (anonymous):

anyone?!?

OpenStudy (anonymous):

you said combine like terms so 6c^2+2c^2=8c^2 @Lyrae

OpenStudy (anonymous):

@LynFran

OpenStudy (anonymous):

yes 8c^2 can you combine the others?

OpenStudy (lynfran):

correct now ..2.5-d is that a typo??

OpenStudy (anonymous):

nope thats what it says so it must be 1d right

OpenStudy (anonymous):

2.5d-1d

OpenStudy (anonymous):

1.5

OpenStudy (lynfran):

2.5d-d-3d=-1.5d

OpenStudy (theedwardsfamily):

First, alll you do is: 6c^2 + 2.5d - d + 2c^2 - 3d combine like terms 6c^2 + 2c^2 = 8c^2 2.5d - 1d - 3d = 2.5d - 4d = - 1.5d then now put them together 8c^2 - 1.5d

OpenStudy (anonymous):

Thanks all of you I wish I could give all of you medals but since @TheEdwardsFamily explained it more she gets the medal.

OpenStudy (theedwardsfamily):

You're welcome! :D

OpenStudy (anonymous):

XD